OpenStudy (anonymous):

If you ride your bike at an average speed of 4km/hour ,and need to travel a total distance of 28km how long will it take you to reach your destination? show work please

OpenStudy (anonymous):

is it like km*T=D 4* 7= 28

OpenStudy (anonymous):

\(\Large v=\frac{d}{t}\) solve for \(t\) \(\Large t=\frac{d}{v}\) where d is the distance, v is the speed and t is the time. so your solution should be something like this: \(\Large t=\frac{28\ km}{4km/hour}\)